A particle strikes a potential barrier of height U and width L. Derive an expression for the approximate transmission probability, if the energy of the particle E < U


1
Expert's answer
2021-05-23T18:55:14-0400

Let's solve the (stationnary) Schrodinger equation in three consecutive domains : before the barrier, in the barries, after the barrier :

  1. "-\frac{\hbar^2}{2m} \psi'' = E \psi , \psi = A_i e^{ikx}+A_re^{-ikx}" , "k = \sqrt{2mE}/\hbar"
  2. "-\frac{\hbar^2}{2m} \psi'' = -(U-E)\psi, \psi = B_i e^{\chi x} + B_r e^{-\chi x}" , "\chi = \sqrt{2m(U-E)}/\hbar"
  3. "-\frac{\hbar^2}{2m} \psi'' = E\psi, \psi = C_i e^{ikx} + C_r e^{-ikx}"

Now by using the continuity of "\psi" and "\psi'", and taking "A_i=1" (as the particle is incident on the barrier) and "C_r=0" (as no particle comes from the other side) we get a set of equations :

"\begin{cases} 1 + A_r = B_i + B_r \\ ik-A_r ik = \chi B_i - \chi B_r \\B_i e^{\chi a}+B_r e^{-\chi a} = C_i e^{ika} \\ B_i \chi e^{\chi a}-\chi B_r e^{-\chi a} = ikC_i e^{ika} \end{cases}"

From the first pair of equations we get "2ik = (\chi +ik)B_i - (\chi-ik)B_r" and thus "B_i = \frac{2ik+(\chi-ik)B_r}{\chi + ik}". We get a system :

"\begin{cases} \frac{2ik}{\chi+ik} e^{\chi a} + B_r(e^{-\chi a}+\frac{\chi - ik}{\chi+ik}e^{\chi a})=C_i e^{ika} \\ \frac{2ik \chi}{\chi+ik} e^{\chi a} - \chi(e^{-\chi a}-\frac{\chi - ik}{\chi+ik}e^{\chi a})B_r=ikC_i e^{ika} \end{cases}"

From this we find "\chi C_i e^{ika}(e^{-\chi a}-\frac{\chi - ik}{\chi+ik}e^{\chi a})+ik C_i e^{ika}(e^{-\chi a}+\frac{\chi - ik}{\chi+ik}e^{\chi a}) = 2\frac{2ik\chi}{\chi+ik}" and thus, finally

"C_i = \frac{4ik\chi e^{-ika}}{(\chi+ik)(\chi e^{-\chi a}-\frac{(\chi-ik)^2}{\chi+ik}e^{\chi a}+ike^{-\chi a})}" , "C_i = \frac{4ik\chi e^{-ika}}{(\chi+ik)^2e^{-\chi a} - (\chi-ik)^2 e^{\chi a}}=\frac{2ik\chi e^{-ika}}{-(\chi^2-k^2)\sinh(\chi a)+2ik\chi \cosh(\chi a)}" , "C_i = \frac{e^{-ika}}{\cosh(\chi a) +i\frac{\chi^2-k^2}{2k\chi}\sinh(\chi a)}"

Thus "T = |C_i|^2 = \frac{1}{\cosh^2(\chi a)+\frac{(\chi^2-k^2)^2}{4k^2\chi^2}\sinh^2(\chi a)} = \frac{1}{1+\frac{(\chi^2+k^2)^2}{4k^2\chi^2}\sinh^2(\chi a)}" , "T=\frac{1}{1+ \frac{((U-E)+E)^2}{4E(U-E)}\sinh^2(\chi a)} = \frac{1}{1+\frac{U^2}{4E(U-E)}\sinh^2(\chi a)}"

When "\frac{U^2}{E(U-E)}\sinh^2(\chi a)\gg 1" (wide, tall barrier) we get "T\sim 16\frac{E}{U}(1-\frac{E}{U})e^{-2\chi a}"